OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 9) and (−1, 11)?

OpenStudy (anonymous):

y + 9 = −(x − 1) y − 9 = 1(x + 1) y + 9 = 1(x + 1) y − 9 = −(x − 1)

OpenStudy (anonymous):

@hartnn @SolomonZelman

hartnn (hartnn):

can you find the slope first ? using the slope formula ? have it ?

OpenStudy (anonymous):

y-9 = (1-x)

hartnn (hartnn):

yes!